# Perturbative soft graviton theorems in de Sitter spacetime

### v1

@@ Abstract
We consider soft graviton scattering for a theory where Einstein’s gravity is minimally coupled to a scalar field in the presence of a cosmological constant, i.e. in a background de Sitter space. Employing a perturbative expansion assuming a small cosmological constant, we compute the leading and subleading corrections to Weinberg’s soft graviton amplitude in the static patch of de Sitter space. We observe similar universal features of the soft graviton amplitude as found in Sayali_Diksha for the soft photons.
@@ 1 Introduction
The infrared properties of gauge theories have been of great interest for a long time, particularly, the soft factorization where the scattering amplitude for processes involving the emission of soft particles is related to the amplitude without soft particles by a universal factor, known as the soft factor, which depends solely on the momentum of the scatterer and the energy of the soft particle Bloch_1937 ; Gell-Mann_Goldberger ; Low_1954 ; Low_1958 ; Weinberg_1965 ; Gross_Jackiw ; Jackiw_1968 ; White_2011 . This soft factorization property of gauge theories, known as soft theorems, later realized as Ward identities of asymptotic symmetries Temple_He ; Laddha_2015 ; Daniel_Kapec ; He_Lysov_Mitra_Strominger ; Lysov_Pasterski_Strominger ; Campiglia_Laddha_2016 ; Kapec_Lysov_2014 ; Campiglia_Laddha_2014 . Another key aspect of infrared physics of gauge theories is the memory effect Zeldovich-Polnarev ; BT ; Strominger_Zhiboedov ; Bieri , which has a trivial mathematical relationship with the soft theorems given by the Fourier transform Strominger_Zhiboedov ; Pasterski_Strominger_Zhiboedov . The memory effect can also be understood as the asymptotic symmetry transformation. The passage of radiation through the future null infinity ($\mathcal{I}^{+}$) produces a transition in the vacuum that is equivalent to the asymptotic symmetry transformation Strominger_Zhiboedov . All such interconnections among three different theories are summarized in the infrared triangle Strominger . There exists many such infrared triangles for the tree-level scattering processes. The loop effects introduce additional terms to the soft factor that depend on the logarithm of the soft energy, resulting in the logarithmic soft theorems Laddha_Sen_2018 ; SS_2019 ; SSS_2020 , which are related to the tail memory effects Laddha_Sen_2019 ; SS_2022 . The asymptotic symmetries corresponding to classical logarithmic soft theorems and tail memory effects have been discovered recently in Choi_Laddha_Puhm_1 , and the complete description of the infrared triangle is given in Choi_Laddha_Puhm_2 ; Choi_Laddha_Puhm_3 .
@@
This paper is organized as follows: In section (2), we briefly describe the geometry of the de Sitter space and define appropriate coordinates for our analysis. We review the scalar fields in the de Sitter background in section (3), which have been studied in details in Sayali_Diksha for generic dimensions $d$. Next, we provide linearized Einstein field equation in the de Sitter background in section (4.1) and obtain mode solutions in section (4.2). By defining the inner product of the modes, we check their orthogonality and consider only the orthogonal set of modes in our analysis. Then, we derive the LSZ formula for gravitons by utilizing the inner product of modes in section (4.3). In section (5), we define the perturbative $\mathcal{S}$-matrix in the static patch of the de Sitter space as in Sayali_Diksha , which differs from the perturbative $\mathcal{S}$-matrix defined in global de Sitter space in Marolf_2013 . Utilizing the LSZ formula, we compute the $\mathcal{S}$-matrix for the scattering process depicted in fig. (2) in section (6). We compute leading and subleading order corrections to the flat space soft graviton theorem in section (6.1) and (6.2), respectively. Finally, we conclude by summarizing our results and providing some remarks in section (7).
@@ 2 De Sitter spacetime
Where, $\eta_{\mu\nu}$ is the Minkowski metric. The advantage of using the above coordinate system is that the metric is conformally flat. We must note that there is a coordinate singularity at $x^{2}=-4l^{2}$, which however does not affect our analysis. Because, we confine a scattering process to a compact region of size $R$ in the static patch (fig. (1)), where $R<<l$. Therefore, all the points in region $R$ have $x^{\mu}<<l$ for every component “$\mu$”, and hence the above metric takes the following form up to $\mathcal{O}(1/l^{2})$:
@@ 3.1 Scalar modes
where $\nabla^{2}$ is the de-Alembertian operator in the de Sitter spacetime and $m$ is the mass of the scalar field $\phi$. Considering corrections up to $\mathcal{O}(1/l^{2})$, this equation has the form
@@ 3.3 Propagator for the scalar field
The scalar field propagator up to $\mathcal{O}(1/l^{2})$ in the de Sitter space has been derived in Sayali_Diksha for arbitrary dimensions $d$. We thus do not run over the full computation, but instead briefly mention the steps required to derive it and then present the final result. The differential equation for the ptopagator can be stated as
@@
Substituting the decomposition of the propagator in terms of $D_{0}$ and $\delta D$ in eq. (16), and performing integration by parts in $p$, the full propagator up to $\mathcal{O}(1/l^{2})$ takes the following form
@@ 3.3.1 Symmetric form of Feynman propagator
Here, we state the symmetric form of Feynman propagator, using the mode solutions $g_{p}$. The Feynman propagator in this case can be decomposed into the flat space propagator and corrections of the order $\mathcal{O}(1/l)$ and $\mathcal{O}(1/l^{2})$, respectively. Thus, we have
@@
As is obvious from the expressions, the propagator stated above is manifestly symmetric in $x$ and $y$. A comparison of this propagator with the asymmetric one is already done in Sayali_Diksha and so we do not re-do it here. We would mention however that $\mathcal{O}(1/l)$ correction $D^{(1)}$ vanishes. This is because we can express the variable $x$ as $-ie^{-ip\cdot x}\partial_{p}(e^{ip\cdot x})$ and convert it into a $y$ dependent piece by doing an integration by parts in $p$. Then, it is not hard to show that these manipulations render $D^{(1)}$ to be zero. We note this is consistent with the absense of a $1/l$ piece in the propagator derived in the previous section. Another thing we want to point out is that although the $1/l^{2}$ correction to the flat space propagator does not seem to match with that in the previous section, we can easily do integration by parts in $p$ and match it thereafter. Thus, all in all, the symmetric form of the Feynman propagator is consistent with its asymmetric form.
@@ 4.1 Linearized Field Equation
where $\Lambda=\dfrac{3}{l^{2}}$ for the de Sitter space. Substituting eqs. (27) and (28) in the above expression, we obtain the linearized form of the Einstein field equation
@@ 4.2 Mode Solutions
In this section, we find the mode expansion for gravitons in the de Sitter spacetime. We first write the linearized field equation (32) in the stereographic coordinates up to $\mathcal{O}(l^{-2})$:
@@
where $\Box=\partial^{\gamma}\partial_{\gamma}$ is the flat space de-Alembertian operator. The above expression is the flat space linearized equation with perturbative corrections up to $\mathcal{O}(l^{-2})$. The de Donder like gauge condition takes the following form
@@ 6 Soft Graviton Theorems
In this section, we compute leading and subleading order corrections to the flat space soft graviton theorem using the LSZ formula discussed in the previous section. We consider a scattering of $n$ scalars with momenta $p_{i}$ followed by an emission of a soft graviton with momentum $k$, as shown in fig. (2). The graviton is attached to an external line by three-point minimal coupling. The graviton attached to an internal line or to higher-point couplings contributes at $\mathcal{O}(k^{0})$, and hence are not relevant to our computation. We consider all external lines to be outgoing for the ease of the calculations. The $\mathcal{S}$-matrix corresponds to the scattering process depicted in fig. (2) is described as
@@
where $G[\phi(z)...\phi(z_{j})]$ is the $n$-point correlation function for scalars, and $\kappa_{i}$ is the scalar-graviton coupling constant. The sum is taken over all the scalars which can emit a soft graviton to take into account the contributions from all the external legs. Now, we perform the Wick contraction among the fields and generate the propagators for the scalars and the graviton.
@@
The above expression can be put in the following form after substituting all essential elements and taking the soft limit, i.e., $k\to 0$.
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\bigg{(}S^{(0)}+\dfrac{1}{l}S^{(1)}+\dfrac{1}{l^{2}}S^{(2)}\bigg{)}\bigg{(}\Gamma_{n}^{(0)}(\{p_{i}\})+\dfrac{1}{l}\Gamma_{n}^{(1)}(\{p_{i}\})+\dfrac{1}{l^{2}}\Gamma_{n}^{(2)}(\{p_{i}\})\bigg{)},$ (55)
where $S^{(0)}$ is the flat space soft factor, $S^{(1)}$ is the leading-order and $S^{(2)}$ is the subleading-order correction to the flat space soft factor. Similarly, $\Gamma_{n}^{(1)}(\{p_{i}\})$ and $\Gamma_{n}^{(2)}(\{p_{i}\})$ are the leading and subleading corrections to the flat space $\mathcal{S}$-matrix, $\Gamma_{n}^{(0)}(\{p_{i}\})$, for the scattering of $n$-scalars. In the following sections, we will substitute the scalar propagator into eq. (54) and compute the soft factors $S^{(0)}$, $S^{(1)}$, $S^{(2)}$.
6.1 Leading corrections to the soft graviton theorem
@@
The above expression contains one zeroth-order $\mathcal{O}(l^{0})$ and three leading-order $\mathcal{O}(l^{-1})$ terms. We analyze them step by step. Let us first evaluate the zeroth-order term:
@@
By performing the $z^{\prime}$-integral in the above expression, one can generate the delta function $\delta^{(4)}(p-p_{i}-k)$, which can be integrated w.r.t. $p$ to replace all $p$ by $p_{i}+k$ in the expression above. Finally, we take the soft limit, i.e. $k\to 0$.
@@
$S^{(0)}=\sum_{i=1}^{n}\kappa_{i}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}+\mathcal{O}(k^{0}).$ (61)
Let us now evaluate the first leading correction term in eq. (57). We can not simply perform the $z^{\prime}$-integral to generate the delta function, because the momentum derivative operator is acting on the exponential. Therefore, we first do the integration by parts in $p$ to move the exponential outside the derivative operator, and then perform the $z^{\prime}$-integral as follows:
@@
where in the last line, the $p$-integral is performed and the soft limit is applied. Similarly, we evaluate the last two leading correction terms in eq. (57). The only difference being that we have to do integration by parts twice to move the exponential outside the derivative operators. After evaluating and collecting all the terms, we obtain
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa_{i}\int d^{4}z\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}-\dfrac{m}{l}\dfrac{(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}+\dfrac{2m}{l}\dfrac{p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\bigg{)}e^{-ip_{i}\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}+\dfrac{1}{l}S^{(0)}\Gamma^{(1)}_{n}(\{p_{i}\})+\mathcal{O}(k^{0})\\ =\sum_{i=1}^{n}\kappa_{i}\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}-\dfrac{im}{l}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}k\cdot\partial_{p_{i}}+\dfrac{2mi}{l}\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\partial_{p_{i}}^{\beta}\bigg{)}\Gamma^{(0)}_{n}(\{p_{i}\})+\dfrac{1}{l}S^{(0)}\Gamma^{(1)}_{n}(\{p_{i}\})+\mathcal{O}(k^{0}).$ (63)
@@
$S^{(1)}=-im\sum_{i=1}^{n}\kappa_{i}\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}k\cdot\partial_{p_{i}}-\dfrac{2p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\partial_{p_{i}}^{\beta}\bigg{)},$ (64)
which is the leading-order correction to the flat space soft factor (61).
@@
  • The gauge transformation does not receive any contributions of $\mathcal{O}(l^{-1})$. Performing the gauge transformation, $\varepsilon_{\alpha\beta}\to\varepsilon_{\alpha\beta}-i\xi_{\alpha}k_{\beta}-i\xi_{\beta}k_{\alpha}$, we have verified that the leading-order correction $S^{(1)}$ to the flat space soft factor is gauge invariant.
@@
  • To check the universal feature of the leading-order corrections, we have used a different set of scalar field modes (13). The new set of modes has no terms of $\mathcal{O}(l^{-1})$, so there are no leading-order corrections to the soft factor for this set of modes. Since the leading-order corrections to the soft factor change for different sets of modes, it is non-universal.
6.2 Subleading corrections to the soft graviton theorem

### v2

@@ Abstract
We consider soft graviton scattering for a theory where Einstein’s gravity is minimally coupled to a scalar field in the presence of a cosmological constant, i.e. in a background de Sitter space. Employing a perturbative expansion in a small cosmological constant, we compute the leading, subleading and sub-subleading corrections to Weinberg’s soft graviton amplitude in the static patch of de Sitter space. We observe similar universal features of the soft graviton amplitude as found in Sayali_Diksha for the soft photons.
@@ 1 Introduction
The infrared properties of gauge theories have been of great interest for a long time, particularly, the soft factorization where the scattering amplitude for processes involving the emission of soft particles is related to the amplitude without soft particles by a universal factor, known as the soft factor, which depends solely on the momentum of the scatterer and the energy of the soft particle Bloch_1937 ; Gell-Mann_Goldberger ; Low_1954 ; Low_1958 ; Weinberg_1965 ; Gross_Jackiw ; Jackiw_1968 ; White_2011 . This soft factorization property of gauge theories, known as soft theorems, later realized as Ward identities of asymptotic symmetries Temple_He ; Laddha_2015 ; Daniel_Kapec ; He_Lysov_Mitra_Strominger ; Lysov_Pasterski_Strominger ; Campiglia_Laddha_2016 ; Kapec_Lysov_2014 ; Campiglia_Laddha_2014 . Another key aspect of infrared physics of gauge theories is the memory effect Zeldovich-Polnarev ; BT ; Strominger_Zhiboedov ; Bieri , which has a trivial mathematical relationship with the soft theorems given by the Fourier transform Strominger_Zhiboedov ; Pasterski_Strominger_Zhiboedov . The memory effect can also be understood as the asymptotic symmetry transformation. The passage of radiation through the future null infinity ($\mathcal{I}^{+}$) produces a transition in the vacuum that is equivalent to the asymptotic symmetry transformation Strominger_Zhiboedov . All such interconnections among three different theories are summarized in the infrared triangle Strominger . There exists many such infrared triangles for the tree-level scattering processes. The loop effects introduce additional terms to the soft factor that depend on the logarithm of the soft energy, resulting in the logarithmic soft theorems Laddha_Sen_2018 ; SS_2019 ; SSS_2020 , which are related to the tail memory effects Laddha_Sen_2019 ; SS_2022 . The asymptotic symmetries corresponding to the logarithmic soft theorems have been discovered recently in Campiglia_Laddha_2019 ; Agrawal_Donnay_2024 ; Choi_Laddha_Puhm_1 , and the complete description of the infrared triangle is given in Choi_Laddha_Puhm_2 ; Choi_Laddha_Puhm_3 .
@@
This paper is organized as follows: In section (2), we briefly describe the geometry of the de Sitter space and define appropriate coordinates for our analysis. We review the scalar fields in the de Sitter background in section (3), which have been studied in details in Sayali_Diksha for generic dimensions $d$. Next, we provide linearized Einstein field equation in the de Sitter background in section (4.1) and obtain mode solutions in section (4.2). By defining the inner product of the modes, we check their orthogonality and consider only the orthogonal set of modes in our analysis. Then, we derive the LSZ formula for gravitons by utilizing the inner product of modes in section (4.3). In section (5), we define the perturbative $\mathcal{S}$-matrix in the static patch of the de Sitter space as in Sayali_Diksha , which differs from the perturbative $\mathcal{S}$-matrix defined in global de Sitter space in Marolf_2013 . Utilizing the LSZ formula, we compute the $\mathcal{S}$-matrix for the scattering process depicted in fig. (2) in section (6). We compute perturbative corrections [^1] By perturbative corrections, we refer to the $\mathcal{O}(l^{-1})$ and $\mathcal{O}(l^{-2})$ corrections to the flat space soft factor, which should not be confused with leading, subleading, and sub-subleading corrections. As analyzed in section (6.2), the $\mathcal{O}(l^{-2})$ corrections include terms of order $1/k^{3}$ and $1/k^{2}$, which do not contribute at the same order. The former is more leading than the latter. Thus, $\mathcal{O}(l^{-2})$ corrections split into subleading and sub-subleading corrections. to the flat space soft factor and deduce leading, subleading, and sub-subleading corrections in the subsequent subsections (6.1) and (6.2). Finally, we conclude by summarizing our results and providing some remarks in section (7).
@@ 2 De Sitter spacetime
Where, $\eta_{\mu\nu}$ is the Minkowski metric. The advantage of using the above coordinate system is that the metric is conformally flat. We must note that there is a coordinate singularity at $x^{2}=-4l^{2}$, which however does not affect our analysis. Because, we confine a scattering process to a compact region of size $R$ in the static patch (fig. (1)), where $R<<l$. Therefore, all the points in region $R$ have $x^{\mu}<<l$ for every component “$\mu$”, and hence the above metric takes the following form up to $\mathcal{O}(l^{-2})$:
@@ 3.1 Scalar modes
where $\nabla^{2}$ is the de-Alembertian operator in the de Sitter spacetime and $m$ is the mass of the scalar field $\phi$. Considering corrections up to $\mathcal{O}(l^{-2})$, this equation has the form
@@ 3.3 Propagator for the scalar field
The scalar field propagator up to $\mathcal{O}(l^{-2})$ in the de Sitter space has been derived in Sayali_Diksha for arbitrary dimensions $d$. We thus do not run over the full computation, but instead briefly mention the steps required to derive it and then present the final result. The differential equation for the ptopagator can be stated as
@@
Substituting the decomposition of the propagator in terms of $D_{0}$ and $\delta D$ in eq. (16), and performing integration by parts in $p$, the full propagator up to $\mathcal{O}(l^{-2})$ takes the following form
@@ 3.3.1 Symmetric form of Feynman propagator
Here, we state the symmetric form of Feynman propagator, using the mode solutions $g_{p}$. The Feynman propagator in this case can be decomposed into the flat space propagator and corrections of the order $\mathcal{O}(l^{-1})$ and $\mathcal{O}(l^{-2})$, respectively. Thus, we have
@@
As is obvious from the expressions, the propagator stated above is manifestly symmetric in $x$ and $y$. A comparison of this propagator with the asymmetric one is already done in Sayali_Diksha and so we do not re-do it here. We would mention however that $\mathcal{O}(l^{-1})$ correction $D^{(1)}$ vanishes. This is because we can express the variable $x$ as $-ie^{-ip\cdot x}\partial_{p}(e^{ip\cdot x})$ and convert it into a $y$ dependent piece by doing an integration by parts in $p$. Then, it is not hard to show that these manipulations render $D^{(1)}$ to be zero. We note this is consistent with the absense of a $1/l$ piece in the propagator derived in the previous section. Another thing we want to point out is that although the $\mathcal{O}(l^{-2})$ correction to the flat space propagator does not seem to match with that in the previous section, we can easily do integration by parts in $p$ and match it thereafter. Thus, all in all, the symmetric form of the Feynman propagator is consistent with its asymmetric form.
@@ 4.1 Linearized Field Equation
where $\Lambda=\dfrac{3}{l^{2}}$ for the de Sitter space. Substituting eqs. (27) and (28) into the above expression, we obtain the linearized form of the Einstein field equation
@@ 4.2 Mode Solutions
In this section, we find the mode expansion for gravitons in the de Sitter space. We first write the linearized field equation (32) in the stereographic coordinates up to $\mathcal{O}(l^{-2})$:
@@
where $\Box=\partial^{\gamma}\partial_{\gamma}$ is the flat space de-Alembertian operator. The above expression is the flat space linearized equation with perturbative corrections up to $\mathcal{O}(l^{-2})$. The de Donder like gauge condition takes the following form in the stereographic coordinates:
@@ 6 Soft Graviton Theorems
In this section, we compute perturbative corrections ($\mathcal{O}(l^{-1})$ and $\mathcal{O}(l^{-2})$ corrections) to the flat space soft graviton theorem using the LSZ formula discussed in the previous section, and deduce leading, subleading and sub-subleading corrections. We consider a scattering of $n$ scalars with momenta $p_{i}$ followed by an emission of a soft graviton with momentum $k$, as shown in fig. (2). The graviton is attached to an external line by three-point minimal coupling. The graviton attached to an internal line or to higher-point couplings are not relevant to our computation. We consider all external lines to be outgoing for the ease of the calculations. The $\mathcal{S}$-matrix corresponds to the scattering process depicted in fig. (2) is described as
@@
where $G[\phi(z)...\phi(z_{j})]$ is the $n$-point correlation function for scalars, and $\kappa_{i}$ is the scalar-graviton coupling constant. The sum is taken over all the scalars which can emit a soft graviton to take into account the contributions from all the external legs. Now, we perform the Wick contraction among the fields and generate the respective propagators.
@@
The above expression can be put in the following form after substituting all essential elements and taking the soft limit, i.e., $\omega\to 0$ (or $k\to 0$).
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},\omega\hat{k})=\bigg{(}A^{(0)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l}A^{(1)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l^{2}}A^{(2)}(\{p_{i}\},\omega\hat{k})\bigg{)}\\ \cdot\bigg{(}\Gamma_{n}^{(0)}(\{p_{i}\})+\dfrac{1}{l}\Gamma_{n}^{(1)}(\{p_{i}\})+\dfrac{1}{l^{2}}\Gamma_{n}^{(2)}(\{p_{i}\})\bigg{)},$ (55)
where $A^{(1)}$ and $A^{(2)}$ are the perturbative corrections to the flat space soft factor $A^{(0)}$. Similarly, $\Gamma_{n}^{(1)}$ and $\Gamma_{n}^{(2)}$ are the perturbative corrections to the flat space $\mathcal{S}$-matrix, $\Gamma_{n}^{(0)}$, for the scattering of $n$-scalars.
We must note that our analysis is valid for $\omega l>1$ in the limits of $l\to\infty$ and $\omega\to 0$. To put the soft factor in a proper hierarchy of leading, subleading, and sub-subleading corrections to the flat space soft factor, one can introduce a small parameter $\delta=1/\omega l$. Thus, the soft factor can be expanded in $\omega$ and $\delta$ as follows:
$A^{(0)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l}A^{(1)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l^{2}}A^{(2)}(\{p_{i}\},\omega\hat{k})=\sum_{r=0}^{2}\sum_{s\geq-1}\delta^{r}\omega^{s}S^{(r,s)}(\{p_{i}\},\hat{k}),$ (56)
where $S^{(0,-1)}$ corresponds to the flat space soft factor, and rest of the coefficients are leading, subleading, sub-subleading corrections, and so on.
6.1 Leading corrections
@@
The above expression contains one $\mathcal{O}(l^{0})$ and three $\mathcal{O}(l^{-1})$ terms. We analyze them step by step. Let us first evaluate the $\mathcal{O}(l^{0})$ term:
@@
By performing the $z^{\prime}$-integral in the above expression, one can generate the delta function $\delta^{(4)}(p-p_{i}-k)$, which can be integrated w.r.t. $p$ to replace all $p$ by $p_{i}+k$ in the expression above. Finally, we take the soft limit, and obtain
@@
$A^{(0)}=\sum_{i=1}^{n}\kappa_{i}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}.$ (62)
One can apply the Ward identity to the above expression and deduce the equivalence principle, i.e. gravity equally couples to all the matter. Thus, $\kappa_{1}=...=\kappa_{n}=\kappa$ [^4] This $\kappa$ should not be confused with $\kappa$ used in section (4.1). Hereafter, $\kappa=\sqrt{8\pi G}$.
$\therefore S^{(0,-1)}=\sum_{i=1}^{n}\kappa\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot\hat{k}}.$ (63)
Let us now evaluate the first $\mathcal{O}(l^{-1})$ term in eq. (58). We can not simply perform the $z^{\prime}$-integral to generate the delta function, because the momentum derivative operator is acting on the exponential. Therefore, we first do the integration by parts in $p$ to move the exponential outside the derivative operator, and then perform the $z^{\prime}$-integral as follows:
@@
where in the last line, the $p$-integral is performed and the soft limit is applied. Similarly, we evaluate the last two $\mathcal{O}(l^{-1})$ terms in eq. (58). The only difference being that we have to do integration by parts twice to move the exponential outside the derivative operators. After evaluating and collecting all the terms, we obtain
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa\int d^{4}z\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}-\dfrac{m}{l}\dfrac{(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}+\dfrac{2m}{l}\dfrac{p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\bigg{)}e^{-ip_{i}\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}+\dfrac{1}{l}A^{(0)}\Gamma^{(1)}_{n}(\{p_{i}\})\\ =\sum_{i=1}^{n}\kappa\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}-\dfrac{im}{l}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}k\cdot\partial_{p_{i}}+\dfrac{2mi}{l}\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\partial_{p_{i}}^{\beta}\bigg{)}\Gamma^{(0)}_{n}(\{p_{i}\})+\dfrac{1}{l}A^{(0)}\Gamma^{(1)}_{n}(\{p_{i}\}).$ (65)
@@
$A^{(1)}=-im\sum_{i=1}^{n}\kappa\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}k\cdot\partial_{p_{i}}-\dfrac{2p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\partial_{p_{i}}^{\beta}\bigg{)}.$ (66)
Thus, the leading correction to the flat space soft factor is
$S^{(1,0)}=-im\sum_{i=1}^{n}\kappa\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot\hat{k})^{2}}\hat{k}\cdot\partial_{p_{i}}-\dfrac{2p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot\hat{k}}\partial_{p_{i}}^{\beta}\bigg{)}.$ (67)
@@
  • The gauge transformation does not receive any contributions of $\mathcal{O}(l^{-1})$. Performing the gauge transformation, $\varepsilon_{\alpha\beta}\to\varepsilon_{\alpha\beta}-i\xi_{\alpha}k_{\beta}-i\xi_{\beta}k_{\alpha}$, we have verified that the leading correction $S^{(1,0)}$ to the flat space soft factor is gauge invariant.
@@
  • To check the universal feature of the leading correction term, we have used a different set of scalar field modes (13). The new set of modes has no terms of $\mathcal{O}(l^{-1})$, so there are no leading corrections to the flat space soft factor for this set of modes. Since the leading corrections to the soft factor change for different set of modes, it is non-universal.
6.2 Subleading and sub-subleading corrections